# Distributional Vision Language Alignment by Cauchy-Schwarz Divergence

### v2

@@ Appendix C Derivations
In CLIP, the features are normalized to compute the loss. Under this unit-norm constraint $\|\mathbf{x}\|_{2}=\|\mathbf{y}\|_{2}=1$, $\mathcal{L}_{\text{InfoNCE}}$ decomposes into
@@
    Dropping the constant $-1/\tau$, define
@@
    Uniformity. For each negative (unpaired) sample $y^{\prime}_{i}$, using Eq. (16),
@@
$\boxed{\mathcal{L}_{\mathrm{clip}}=\mathcal{L}_{\mathrm{align}}+\mathcal{L}_{\mathrm{uniform}}+\text{const.}}$.
@@ D.1 Connection to the prior loss
Connection to the prior loss ($\ell_{2}$ loss) used by Eclipse [15]:
@@
$D_{\mathrm{KL}}(p;q)=\int p(\omega)\log\frac{p(\omega)}{q(\omega)}\,d\omega$ (21)
@@
[…]
@@
Nonparametric estimation. Additionally, when the distributions are not assumed to be Gaussian, a nonparametric estimator is required for KL divergence. A common choice, the k-NN estimator [52], is non-differentiable, which poses challenges for optimization in gradient-based learning frameworks. In contrast, the CS divergence demonstrates greater stability and differentiability when paired with KDE, making it a more robust choice.
@@
Wasserstein Distance is also widely used for distribution discrepancy (e.g. GAN [53]). However, Wasserstein distance is be computed either by using an additional learnable module (e.g., a neural network for estimating a transport map [54]) or by solving an optimization problem, often approximated via multiple Sinkhorn [55] iterations for computational efficiency, leading to efficiency problem in large-scale training. In contrast, CS divergence can be efficiently estimated by a nonparametric estimator.
D.4 Comparison with mutual information divergence [56].
Mutual information estimation depends on parametric assumptions about the underlying distributions—e.g., multivariate Gaussian—whereas CS divergence imposes no such constraints. Moreover, estimating mutual information decomposes into a mutual information term plus two KL divergences, and thus lacks explicit connections to the InfoNCE in terms of alignment and uniformity.
Appendix E More Results
E.1 More Visualization
@@
E.2 More visualization for token alignment
@@
Appendix F Implementation details
@@
Our models were trained on 4 NVIDIA RTX A100 GPUs with a global batch size of 1,024 (256 per GPU). We optimized parameters using AdamW with a cosine annealing learning rate schedule, spanning a total of 100 GPU hours. Mixed-precision training (FP16) was employed to enhance computational efficiency while maintaining stability. We use the learning rate of $55e-$ 5. We use hyperparameter $\lambda$ as $0.01$ to keep the same number scale as the divergence.
@@
A proper kernel size is critical in KDE for accurate estimation of Eq. (8). In this paper, we follow [51] to normalize the features from two modalities and use a kernel size $1$. In general, this is sufficient to ensure stable learning.
F.1 T2I details
@@
We use Kandinsky v2.2, an unCLIP-type model that utilizes CLIP ViT-bigG-14-laion2B-39B-b160k with 1280 projection dimensions for text and image encoders. Kandinsky v2.2 employs a latent diffusion model and MOVQ [57] as the decoder to generate images of size $512\times 512$ from the given image representation. When using the Kandinsky decoder, we apply $50$ denoising steps [58] with a classifier-free guidance scale of 7.5 [59].
@@
To ensure a fair comparison, our adapter module has the same architecture as Eclipse [15], which is based on the standard PriorTransformer model [3] but modified to be time-independent. Specifically, it consists of 10 layers with 16 attention heads, each having a head dimension of 32. The embedding dimension is 768/1280, with three additional embeddings. The model does not use time embeddings and has a dropout rate of 0.0. For the text to image generation task, in order to use the pretrained image generator, we only use the text adapter. For the retrieval and classification, we use adapters for both modalities.
@@
Appendix G More experimental results
@@
We compare with CLIP-Adapter [13] on the image classification task following their few-shot classification setting. We fine-tune the adapter based on ViT-B/16 with 16-shots subset for each of the 11 datasets. The results are provided in the following table. With better alignment, our method consistently performs better.
@@
Table 6 shows that our method is robust to moderate variations in both $\lambda$ and $\sigma$, with only minor FID fluctuations over a wide range.
@@
We train the adapter solely with InfoNCE and use the Kandinsky decoder to generate the corresponding images. Table 7 shows that InfoNCE alone struggles to align the multimodal distributions, resulting in a high FID score.
Table 7: Ablation study of CS-Aligner. Alignment with CS-Aligner significantly outperforms the alignment results with only InfoNCE.
Loss FID
InfoNCE 151.35
CS-Aligner 12.62

### v3

@@ Appendix B Derivations
In CLIP, the features are normalized to compute the loss. Under this unit‐norm constraint $\|\mathbf{x}\|_{2}=\|\mathbf{y}\|_{2}=1$, $\mathcal{L}_{\text{InfoNCE}}$ decomposes into
@@
    Dropping the constant $1/\tau$, define
@@
    Uniformity. For each negative (unpaired) sample $y^{\prime}_{i}$, using Eq. equation 16,
@@
$\boxed{\mathcal{L}_{\mathrm{clip}}=\mathcal{L}_{\mathrm{align}}+\mathcal{L}_{\mathrm{uniform}}+\text{const.}}$
Appendix C Related Work of Cauchy-Schwarz (CS) divergence.
CS divergence (Principe et al., 2000a; b) is derived from the Cauchy-Schwarz inequality for square-integrable functions. It serves as a symmetric distribution distance metric with notable properties, such as the ability to measure conditional distributions (Yu et al., 2025) and the closed-form expression for mixtures of Gaussians (Kampa et al., 2011). CS divergence has been successfully applied across various domains, including deep clustering (Trosten et al., 2021), disentangled representation learning (Tran et al., 2022), and deep regression (Yu et al., 2024). Moreover, due to its advantage of estimating discrepancy between conditional distributions, it has demonstrated success in the domain adaption area (Yin et al., 2024) and time series clustering (Yu et al., 2025). However, the utility of CS divergence in foundation models remains unclear and unexplored.
@@ D.1 Connection to the prior loss
Connection to the prior loss ($\ell_{2}$ loss) used by Eclipse (Patel et al., 2024):
@@
$D_{\mathrm{KL}}(p;q)=\int p(\omega)\log\frac{p(\omega)}{q(\omega)}\,d\omega.$ (21)
@@
In contrast, the CS divergence becomes infinite only if there is no overlap between supports of $p$ and $q$, i.e., when $\int p(\omega)q(\omega)d\omega=0$, making the logarithm undefined. Hence, the condition for finite CS divergence is: $\text{supp}(p)\cap\text{supp}(q)\neq\emptyset.$
@@
Nonparametric estimation. Additionally, when the distributions are not assumed to be Gaussian, a nonparametric estimator is required for KL divergence. A common choice, the k-NN estimator (Wang et al., 2009), is non-differentiable, which poses challenges for optimization in gradient-based learning frameworks. In contrast, the CS divergence demonstrates greater stability and differentiability when paired with KDE, making it a more robust choice.
@@
Wasserstein Distance is also widely used for distribution discrepancy (e.g. GAN (Arjovsky et al., 2017)). However, Wasserstein distance is be computed either by using an additional learnable module (e.g., a neural network for estimating a transport map (Korotin et al., 2022)) or by solving an optimization problem, often approximated via multiple Sinkhorn (Cuturi, 2013) iterations for computational efficiency, leading to efficiency problem in large-scale training. In contrast, CS divergence can be efficiently estimated by a nonparametric estimator.
D.4 Quantitative comparisons with KL and Wasserstein distance.
We compare our method with KL and Wasserstein distances below. To make the KL divergence tractable, we assume the batch embeddings follow Gaussian distributions. For the Wasserstein distance, we either use the closed-form Gaussian Wasserstein distance under the same assumption or apply the Sinkhorn algorithm for general distributions. However, in practice, we found that Sinkhorn often fails to converge. The results show that our method outperforms both KL and Wasserstein distances. Moreover, Wasserstein distance and KL lack an InfoNCE-style alignment–uniformity decomposition; only CS-divergence yields the compatible formulation (Eq. 10). The Gaussian assumption is also stronger than our nonparametric method.
Method FID $\downarrow$
KL 23.48
W-distance 18.41
Sinkhorn Not converge
CS-Aligner 12.62
D.5 Comparison with mutual information divergence (Kim et al., 2022).
Mutual information estimation depends on parametric assumptions about the underlying distributions, e.g., multivariate Gaussian, whereas CS divergence imposes no such constraints. Moreover, estimating mutual information decomposes into a mutual information term plus two KL divergences, and thus lacks explicit connections to the InfoNCE in terms of alignment and uniformity.
Appendix E Computation complexity and stability analysis
We normalize high-dimensional embeddings onto the unit hypersphere and use a fixed Gaussian kernel bandwidth so that concentration of measure and classical KDE theory ensure stable, low-variance estimates.
In high dimensions, mapping embeddings onto the unit hypersphere exploits the concentration of measure phenomenon: as $d$ grows, the pairwise distances $\|x-y\|$ between random points on $S^{d-1}$ concentrate sharply around $\sqrt{2}$, with fluctuations of order $O(1/\sqrt{d})$. Consequently, a Gaussian kernel
$K(x,y)=\exp\!\Bigl(-\frac{\|x-y\|^{2}}{2\sigma^{2}}\Bigr),$ (24)
with fixed bandwidth (e.g. $\sigma=1$) yields values confined to a narrow, well-behaved range, preventing weights from collapsing to 0 or saturating at 1 and ensuring smoothly varying density estimates (Berestycki and Nickl, 2009).
Moreover, when the effective sample size $n$ (e.g. batch size) and dimensionality $d$ satisfy $n\,\sigma^{d}\;\gg\;1$, which holds for $\sigma=1$, $n\sim 10^{3}$, and $d\sim 10^{3}$, the KDE estimator obeys a central limit theorem. This guarantees that CS divergence estimates have vanishing variance and stable gradients during optimization (Parzen, 1962).
Computational complexity.
The computation cost of our method is comparable to the CLIP-based method when scaling up to even larger-scale datasets. The computation complexity is $O(N^{2})$, which is the same as the InfoNCE used in CLIP. However, the computational complexity is feasible to scale up to larger-scale datasets.
Appendix F More Results
F.1 More Visualization
@@
F.2 More visualization for token alignment
@@
Appendix G Implementation details
@@
Our models were trained on 4 NVIDIA RTX A100 GPUs with a global batch size of 1,024 (256 per GPU). We optimized parameters using AdamW with a cosine annealing learning rate schedule, spanning a total of 100 GPU hours. Mixed-precision training (FP16) was employed to enhance computational efficiency while maintaining stability. We use the learning rate of 5e55e-5. We use hyperparameter $\lambda$ as $0.01$ to keep the same number scale as the divergence.
@@
A proper kernel size is critical in KDE for accurate estimation of Eq. (8). In this paper, we follow Yin et al. (2024) to normalize the features from two modalities and use a kernel size 11. In general, this is sufficient to ensure stable learning.
G.1 T2I details
Figure 1 implementation details.
For Fig. 1 and Fig. 1, we train the same adapter on top of the CLIP model using InfoNCE and CS-Aligner, respectively. We use the MSCOCO training set and visualize the learned representations with t-SNE on 5K image–text pairs from the validation set. For the temperature in both InfoNCE and CS-Aligner, we initialize it from the pretrained CLIP model and keep it learnable during training. For Fig. 1, we compute the L2 distance between the embeddings of all image–text pairs and visualize the resulting histogram. The histogram of L2 distances for positive pairs systematically reflects their distance distribution.
@@ G.1 T2I details
We use Kandinsky v2.2, an unCLIP-type model that utilizes CLIP ViT-bigG-14-laion2B-39B-b160k with 1280 projection dimensions for text and image encoders. Kandinsky v2.2 employs a latent diffusion model and MOVQ (Zheng et al., 2022) as the decoder to generate images of size $512\times 512$ from the given image representation. When using the Kandinsky decoder, we apply $50$ denoising steps (Ho et al., 2020) with a classifier-free guidance scale of 7.5 (Ho and Salimans, 2022).
@@
To ensure a fair comparison, our adapter module has the same architecture as Eclipse (Patel et al., 2024), which is based on the standard PriorTransformer model (Ramesh et al., 2022) but modified to be time-independent. Specifically, it consists of 10 layers with 16 attention heads, each having a head dimension of 32. The embedding dimension is 768/1280, with three additional embeddings. The model does not use time embeddings and has a dropout rate of 0.0. For the text to image generation task, in order to use the pretrained image generator, we only use the text adapter. For the retrieval and classification, we use adapters for both modalities.
@@
Appendix H More experimental results
@@
We compare with CLIP-Adapter (Gao et al., 2024) on the image classification task following their few-shot classification setting. We fine-tune the adapter based on ViT-B/16 with 16-shots subset for each of the 11 datasets. The results are provided in the following table. With better alignment, our method consistently performs better.
@@
Table 6: Image captioning results.
Method Bleu_1 $\uparrow$ CIDEr $\uparrow$
InfoNCE+LM 40.4 14.3
InfoNCE+LM+CS 41.3 16.7
Image captioning.
We extend our method to the image captioning task. We adopt the Blip2 (Li et al., 2023a) stage one training strategy to highlight the importance of representation alignment for the image captioning task. We train a Q-former with the image text contrastive loss (InfoNCE) and the language model loss on the MSCOCO captioning dataset. The results in Table 6 show that our method can improve the image captioning ability. Also, the qualitative results of image captioning are shown in Fig. reffig:vis-captioning. The generated captions are semantically aligned with the images, demonstrating the general applicability of our method.
Figure 8: Qualitative results of image captioning.
@@
H.1 Ablation study
Table 7 shows that our method is robust to moderate variations in both $\lambda$ and $\sigma$, with only minor FID fluctuations over a wide range. A large $\lambda$ overemphasizes distributional alignment, optimizing intra-modality uniformity and global distribution distance while overlooking the pairwise alignment term. Since the generation task is sensitive to both global distribution closeness and sample-wise alignment, an excessively large $\lambda$ can degrade performance. We also evaluate the sensitivity of $\lambda$ and $\sigma$ on the MSCOCO retrieval task. The results show that our method is robust and performs well across a wide range of hyperparameters.
Table 8: Sensitivity to $\lambda$ and $\sigma$. on MSCOCO retrieval
$\sigma$ 0.1 0.5 1 1.5
R@1 49.3 50.9 50.7 50.2
$\lambda$ 0.01 0.1 1 10
R@1 50.1 50.6 50.7 48.6
@@ H.1 Ablation study
We ablate InfoNCE and InfoNCE with CS divergence (CS-Aligner) on the text-to-image generation task. Specifically, we train the adapter on the MSCOCO dataset and use the Kandinsky decoder to generate the corresponding images. For the InfoNCE temperature, we resume it from the pretrained CLIP model and keep it learnable. We then compute the FID score for comparison (lower is better). Table 9 shows that InfoNCE alone struggles to align the multimodal distributions, resulting in a high FID score. As the learnable temperature $\tau$ (inherited from CLIP) decreases during training, the contrastive logits become sharper, making the uniformity term dominate over the alignment term and thereby weakening multimodal alignment (see our decomposition in Sec. 2) For text-to-image generation, the decoder requires the two modalities to lie in the same distribution, which InfoNCE alone is unable to guarantee. Hence, an InfoNCE-only model may still perform well in cosine-similarity–based retrieval but fails in generation due to the persistent distributional gap.
We also provide the retrieval ablations. Retrieval requires only correct relative similarity ranking, not full distributional overlap, so the degradation of InfoNCE-only is smaller. Nevertheless, CS-Aligner consistently outperforms InfoNCE-only, likely because the intra-modality uniformity terms (Eq. 9) promote better sample separability, which benefits retrieval.
Table 9: Ablation study of CS-Aligner on retrieval and generation. Alignment with CS-Aligner significantly outperforms using InfoNCE alone.
Method Retrieval Generation
T2I I2T Avg FID $\downarrow$
InfoNCE 50.1 65.8 57.95 151.35
CS-Aligner 50.7 66.34 58.52 12.62